Show that

a)z+z*=2 Re z=2x

b)z-z*=2i Im z=2iy

c)z/z*={x^2-y^2/x^2+y^2}+i{2xy/x^2+y^2}


1
Expert's answer
2020-08-23T17:58:25-0400

"z=x+iy\newline
z^*=x-iy\newline
a) z+z^*=x+iy+x-iy=2x\newline
b) z-z^*=x+iy-(x-iy)=2iy\newline
c) \dfrac{z}{z^*}=\dfrac{x+iy}{x-iy}=\dfrac{(x+iy)*(x-iy)}{(x-iy)*(x-iy)}=\dfrac{x^2+2ixy+y^2}{x^2-y^2}=\dfrac{x^2+y^2}{x^2-y^2}+\dfrac{2ixy}{x^2-y^2}"